OpenStudy (anonymous):

If b = 3, then b^-2 is equivalent to _____. 9 -9 -6 1 over 9

OpenStudy (anonymous):

what do you mean by b^2?

OpenStudy (ajprincess):

b^-2=1/b^2

OpenStudy (anonymous):

B is 3

OpenStudy (anonymous):

its means 3^-2

OpenStudy (ajprincess):

ya. 3^-2=1/3^2

OpenStudy (ajprincess):

3^2=3*3

OpenStudy (ajprincess):

Can u find it nw?

OpenStudy (anonymous):

Its 9?:D

OpenStudy (ajprincess):

ya 3^2=9 bt vat is 3^-2?